A 200 g ball is shot up a vertical

column of viscous liquid with an

initial velocity of 10 m/s Assuming

the force impeding the ball's motion

because of the viscosity of the liquid

is constant at 4.5 N, calculate how

high the ball goes before coming to

rest.



1
Expert's answer
2021-01-31T19:26:38-0500

When a ball is hit upwards the maximum height of a projectile is given by


"H=u^2y\/2g"

"u"o "y" ="u"o "sin""\theta"

mass=200"g"

=0.2kg

innitial velocity("u)" ="10m\/s"

in this case, we can get the normal force acting on a body vertically due to gravity.

"f"m="mg"

="0.2*9.8"

="1.96N"

viscosity of the liquid="4.5N"

"tan\theta" "=4.5\/1.96"

"tan" -1=2.2959

"\theta" "=66.46"0


from "h"max"=" "u^2*sin(\theta"")^2" "\/2g"


"=" "10^2*sin(66.46)^2" "\/2*9.8"


"=4.2884m"